OpenStudy (anonymous):

3x + 2y = 4 2x - y = 5 This system of equations

OpenStudy (anonymous):

A)has no solution. B) has one solution C) is coincident

OpenStudy (solomonzelman):

3x + 2y = 4 2x - y = 5 --> times 2 --> 4x - 2y = 10 so you now have 3x + 2y = 4 4x - 2y = 10 ------------- 7x + 0y = 14 (I added them to each other) 7x=14 x=2 Solve for y. 2x - y = 5 plug in 2 for x 2(2)- y = 5 4-y=5 4-y-4=5-4 -y=1 y=-1 check it using the first equation since the second we already know, 3(2) + 2(-1) = 4 6-2=4 4=4 So x=2 y=-1 OR (2,-1)